# Compressing neural network by tensor network with exponentially fewer variational parameters

### v2

@@ III Compression performance
The approximation holds since in general we have #$\#(\text{ADTN})\ll\#(\boldsymbol{T}^{\text{res}})\ll\#(\text{NN})$. Therefore, $\rho_{\text{tot}}$ can be significantly improved by simply compressing more NN layers to ADTN’s. Our results suggest that for $\rho\lesssim 0.3$, shallow ADTN (say with $M=1$ tensor layer) is sufficient for compressing the chosen parts of the NN. The training accuracy $\tilde{\eta}$ is almost $100\%$, meaning the representational ability of the shallow ADTN is sufficient. As $\rho$ decreases by compressing more NN layers, $\tilde{\eta}$ also decreases. Deeper ADTN with better representational ability is required. As shown in the last row, the testing accuracy is improved from $72.13\%$ to $78.95\%$ by increasing $M$ from $3$ to $9$ when compressing 12 layers in VGG-16. The total compression ratio is lowered to $\rho_{\text{tot}}\simeq 0.008$.
IV Discussions on over-parametrization, compression order, and faithfulness
@@
Below, we further discuss on several critical issues about the ADTN compression scheme. First, we show that by properly increasing the number of ADTN’s ($N$), the NN can be well compressed whether it is under-fitting or over-fitting. We take LeNet-5 as an example and alter the output dimension of the first linear layer $s$ (i.e., the input dimension of the second linear layer) to change the parameter complexity of the NN. Fig. 2 (a) shows the testing-accuracy ratio $\eta/\eta_{\text{NN}}$ versus the inverse total compression ratio $\rho_{\text{tot}}^{-1}$ for the CIFAR-10 dataset with $s$ ranging from $32$ to $1536$.
@@
The NN with a small $s$ is difficult to compress even by ADTN. In this case, the model is under-fitting, resulting in small $\rho_{\text{tot}}^{-1}$ and $\eta/\eta_{\text{NN}}$ (see the symbols with relatively light colors). By increasing $N$, $\eta/\eta_{\text{NN}}$ will fairly approach to $1$. In comparison, for the over-fitting case with large $s$, the NN is shown to be compressible by ADTN, leading to large $\rho_{\text{tot}}^{-1}$ and $\eta/\eta_{\text{NN}}>1$, particularly when using multiple ADTN’s for compression ($N>1$). See the symbols with dark colors. The testing accuracy is increased by compression with $\eta/\eta_{\text{NN}}>1$, implying that the generalization ability is improved. Generally speaking, our results suggest that compressing by multiple ADTN’s will avoid significant decrease of the testing accuracy in the under-fitting cases, and enhance generalization ability in the over-fitting cases.
@@
Figure 2: (Color online) The testing-accuracy ratio $\eta/\eta_{\text{NN}}$ versus the inverse total compression ratio $\rho_{\text{tot}}^{-1}$ for the CIFAR-10 dataset by LeNet-5. The dimensions of the first two linear layers are taken as (512×(512$\times$( 512 ×s)))) and $(s\times 128)$, where $s$ is taken as $32,64,\cdots,1536$ (see the color bar). $N$ ADTN’s are used for compression, where each contains $M=3$ tensor layers. The testing-accuracy ratio $\eta/\eta_{\text{NN}}$ versus the inverse of compression ratio $\rho^{-1}$ for the CIFAR-10 dataset by VGG-16. We selected $12$ largest layers (ten convolutional and two linear layers) in VGG-16 and compress them in two different orders: from the input to output (denoted as forward compression) and the other way around (denoted as backward compression).
@@
Considering compressing multiple layers in a NN, one will suffer from severe local-minima problem. Simultaneously compressing several NN layers will be unstable since the optimization can easily be trapped in different local minima when taking different initializations of the ADTN’s. A stable way is to compress layer by layer. Specifically, we start from the compression of one layer by encoding it as ADTN(s). After this is done, we compress another layer by optimizing the ADTN(s), and simultaneously optimize the ADTN(s) of the formerly-compressed layer. For each time the optimization process converges, we consider to compress one more layer by optimizing the ADTN(s) of this layer and those of all optimized layers simultaneously. This requires to determine a compression order.
@@
Below we show that the compression order should be “backward”, meaning compressing from the layers closer to the output of NN to those nearer the input. Fig. 2 (b) shows the $\eta/\eta_{\text{NN}}$ versus $\rho^{-1}$ for the CIFAR-10 dataset by the reduced VGG-16. We compress 12 largest layers therein from the input to output (forward) and the other way around. Much better performance is achieved by the backward compression. This is essentially due to the interdependence among the compressions of different layers. An accurate compression of a NN layer should require accurate input from the former layers in the information propagation process. We conjecture the above observation to be held when using other compression means.
@@
We expect a well-performed compression scheme to at least faithfully restore the generalization ability of the original NN, despite the NN itself is well trained or not. It means that the testing accuracy after compression will be mainly determined by the testing accuracy of the original NN. Fig. 3 shows the testing accuracies before and after compression (denoted as $\eta_{\text{NN}}$ and $\eta$) for VGG-16 on the CIFAR-10 dataset. Different before-compression testing accuracies are obtained by changing the number of training samples. Our results show that the ADTN compression scheme faithfully restores the testing accuracies of the NN with slight improvements. We expect the ADTN scheme to be faithful for compressing other NN models.
V Summary
In summary, we proposed a general compression scheme based on automatically differentiable tensor network (ADTN) to significantly reduce the variational parameters of neural networks (NN’s). The key idea is to encode the higher-order parameter tensors of NN as the contractions of deep ADTN(s) that contain exponentially less parameters. The performance of our scheme has been demonstrated with several well-known NN’s on different datasets. Discussions on over-parametrization, compression order, and faithfulness of our ADTN compression scheme are given. Our work suggests TN as a more efficient and compact mathematical form than the multi-way arrays to represent the variational parameters of NN’s.
@@ Acknowledgments
YQ is thankful to Ze-feng Gao for stimulating discussions. This work was supported by NSFC (Grant No. 12004266), Beijing Natural Science Foundation (Grant No. 1232025), Tianjin Natural Science Foundation of China (Grant No. 20JCYBJC00500), and Academy for Multidisciplinary Studies, Capital Normal University.
Appendix A Basics of tensor network and the contractions
@@
Appendix B Details of FC-2, LeNet-5, AlexNet, AFNet and VGG-16

### v3

@@ III.1 Compression Performance
Our results suggest that for $\rho\lesssim 0.3$, shallow ADTN (say with $M=1$ tensor layer) is sufficient for compressing the chosen parts of the NN. The training accuracy $\tilde{\eta}$ is almost $100\%$, meaning the representational ability of the shallow ADTN is sufficient. As $\rho$ decreases by compressing more NN layers, $\tilde{\eta}$ also decreases. Deeper ADTN with better representational ability is required. As shown in the last row, the testing accuracy is improved from $72.13\%$ to $78.95\%$ by increasing $M$ from $3$ to $9$ when compressing 12 layers in VGG-16. The total compression ratio is lowered to $\rho_{\text{tot}}\simeq 0.008$.
Table 2 compare the performances of ADTN with various schemes, including Tucker decomposition, CPD, MUSCO, and MTD. We take the data reported in Refs. [[25]] and [[26]] for comparison. Three typical results of ADTN are given in the lower half of Table 2. The mark ($N$-$M$) of ADTN indicates using $N$ ADTN’s, each containing $M$ TN layers, for compression. $\mathcal{M}_{\text{NN}}$ is the number of compressed NN layers, with totally 16 layers in VGG-16. ADTN demonstrates remark compression performances. With ($6$-$3$), our ADTN scheme compressing 8 NN layers achieves the highest testing accuracy $\eta=91.90\%$, with the total compressions ratio $\rho_{tot}=1/17.65$. This clearly surpasses Tucker decomposition and CPD. Additionally, with ($6$-$3$) for compressing 11 NN layers, much better total compression ratio $\rho_{\text{tot}}=1/73.47$ is achieved, which far surpasses the total compression ratios obtained by the existing methods.
VGG-16 $\mathcal{M}_{\text{NN}}$ $1/\rho_{\text{tot}}$ $\eta$ $\eta/\eta_{NN}$
Baseline [26] - - 90.63% -
Tucker [26] - 11.82 91.02% 100.43%
CPD [26] - 11.76 87.02% 96.02%
MUSCO [25] - 23.55 87.07% 96.07%
MTD [26] - 32.59 89.80% 99.08%
Baseline (our) - - 90.54% -
ADTN (2-3) 10 46.97 89.85% 99.24%
ADTN (6-3) 8 17.56 91.90% 101.50%
ADTN (6-3) 11 73.47 88.55% 97.80%
Table 2: The inverse total compression ratios $1/\rho_{\text{tot}}$, after-compression testing accuracy $\eta$, and the ratio against the baseline accuracy with various existing compression schemes and ADTN. $\mathcal{M}$ denotes the number of compressed NN layers by ADTN. The mark ($N$-$M$) specifies the number of ADTN’s $N$ used for compressing one NN layer and the number of TN layers $M$ in each ADTN. The compressed NN in this table is VGG-16 by considering the CIFAR-10 dataset.
We also compare ADTN with the compression scheme based on MPO [34], another typical shallow TN. MPO has a similar structure to MPS (TT form), and is used to represent quantum operators instead of states. Fig. 3 shows the testing accuracy $\eta$ on CIFAR-10 after compressing the last convolutional layers in VGG-16 versus the inverse compression ratio $1/\rho$. The number of parameters of MPO is adjusted by varying its virtual bond dimensions. The number of TN layers in the ADTN ranges from $M=1$ to $5$. Both methods exceeds the baseline accuracy (horizontal dash line), while ADTN generally achieves higher inverse compression ratio. This is in accordance with previous results in the relevant fields, which will be discussed in Sec. 4.4.
III.2 Discussions on compression with under/overfitting
Figure 3: (Color online) The after-compression testing accuracy $\eta$ on CIFAR-10 versus the inverse compression ratio $1/\rho$, obtained by compressing a simplified VGG-16 using ADTN and MPO (shallow TN). The baseline testing accuracy is $\eta_{\text{NN}}=0.8114$.
@@ III.2 Discussions on compression with under/overfitting
Below, we further discuss on several critical issues about the ADTN compression scheme. First, we demonstrate that by properly increasing the number of ADTN’s ($N$), the NN can be effectively compressed whether it is under-fitting or overfitting. We take LeNet-5 as an example and alter the output dimension of the first linear layer $s$ (i.e., the input dimension of the second linear layer) to change the parameter complexity of the NN. Fig. 4 (a) shows the testing-accuracy ratio $\eta/\eta_{\text{NN}}$ versus the inverse total compression ratio $\rho_{\text{tot}}^{-1}$ for the CIFAR-10 dataset with $s$ ranging from $32$ to $1536$.
@@
The NN with a small $s$ is difficult to compress even by ADTN. In this case, the model is under-fitting, resulting in small $\rho_{\text{tot}}^{-1}$ and $\eta/\eta_{\text{NN}}$ (see the symbols with relatively light colors). By increasing $N$, $\eta/\eta_{\text{NN}}$ will fairly approach to $1$, mitigating the harm of accuracy.
In contrast, for large $s$ where the NN is overfitting, the NN is shown to be compressible by ADTN, leading to large $\rho_{\text{tot}}^{-1}$ and $\eta/\eta_{\text{NN}}>1$, particularly when using multiple ADTN’s for compression ($N>1$). See the symbols with dark colors. The testing accuracy is increased by compression with $\eta/\eta_{\text{NN}}>1$, implying that the generalization ability is improved. Generally, our results suggest that compressing by multiple ADTN’s will avoid significant decrease of the testing accuracy in the under-fitting cases, and enhance generalization ability in the overfitting cases.
@@
Figure 4: (Color online) (a) The testing-accuracy ratio $\eta/\eta_{\text{NN}}$ versus the inverse total compression ratio $\rho_{\text{tot}}^{-1}$ for the CIFAR-10 dataset by LeNet-5. The dimensions of the first two linear layers are taken as $(512\times s)$ and $(s\times 128)$, where $s$ is taken as $32,64,\cdots,1536$ (see the color bar). $N$ ADTN’s are used for compression, where each contains $M=3$ TN layers. (b) The testing-accuracy ratio $\eta/\eta_{\text{NN}}$ versus the inverse of compression ratio $\rho^{-1}$ for the CIFAR-10 dataset by VGG-16. We selected $12$ largest layers (ten convolutional and two linear layers) in VGG-16 and compress them in two different orders: from the input to output (denoted as forward compression) and the other way around (denoted as backward compression).
III.3 Discussions on local-minima problem and compression order
@@
Considering compressing multiple layers in an NN, one suffers from severe local-minima problem. Simultaneously compressing several NN layers can be unstable, as optimization process may easily be trapped in different local minima when taking different initializations of the ADTN’s. A stable way we have found is to compress layer by layer.
Specifically, we start from the compression of one layer by encoding it as ADTN(s). After this is done, we compress another layer by optimizing the ADTN(s), and simultaneously optimize the ADTN(s) of the formerly compressed layer. For each time the optimization process converges, we consider compressing one more layer by optimizing the ADTN(s) of this layer and those of all optimized layers simultaneously. This requires determining a compression order.
@@
We show that the compression order should be “backward”, meaning compressing from the layers closer to the output of NN to those nearer the input. Fig. 4 (b) shows the $\eta/\eta_{\text{NN}}$ versus $\rho^{-1}$ for the CIFAR-10 dataset by VGG-16. We compress 12 largest layers therein from the input to output (forward) and the other way around (backward). Significantly better performance is achieved with the backward compression. This is essentially due to the interdependence among the compressions of different layers. An accurate compression of a NN layer should require accurate input from the former layers in the information propagation process. We conjecture the above observation to be held when using other compression means.
III.4 Discussions on Compression Faithfulness
We expect a well-performing compression scheme to at least faithfully restore the generalization ability of the original NN, regardless of whether the NN itself is well-trained or not. This means the testing accuracy after compression will be (at least) close to the testing accuracy of the original NN (the baseline). Fig. 5 shows the testing accuracies before and after compression (denoted as $\eta_{\text{NN}}$ and $\eta$) for VGG-16 on the CIFAR-10 dataset. Different testing accuracies before compression are obtained by varying the number of training samples. Our results show that the ADTN compression scheme faithfully restores the testing accuracies of the NN with slight improvements. We expect the ADTN scheme to be faithful for compressing other NN models as well.
@@ III.4 Discussions on Compression Faithfulness
III.5 Discussions on Tensor Network: “deep” versus “shallow”
Deep and shallow TN’s possess essential differences in the fields of quantum physics, quantum computation, and quantum-inspired machine learning. Here, we specifically refer to the shallow TN’s as those with just one layer. In quantum many-body physics, MPS (a shallow TN) can well approximate the gapped states of one-dimensional (1D) quantum systems [46], where the entanglement entropy satisfies the so-called 1D area law. The projected entangled pair state (PEPS), which belongs to deep TN’s, can approximate the states with much higher entanglement (obeying the two-dimensional area law) [47, 48].
Note entanglement closely associates with the nonsparsity or more specifically, the number of nonvanishing singular values and thus closely associates with the compressibility. In other words, deep TN’s can generally approximate quantum states whose coefficients form much denser tensors compared to those approximated by shallow TN’s. Sparsity is closely related to compressibility. A tensor, whether it stores some data or variational parameters, possesses certain compressibility if it is sparse. This coincides with our results, where deep ADTN exhibits higher compression performance over the schemes based on MPS (TT form) and MPO. Similar observations have been made in the TN-based quantum-inspired machine learning [49] where PEPS outperforms MPS for classification and generation [50].
A more explicit example demonstrating the superior power of deep TN’s over the shallow ones is the preparation of MPS by quantum circuit. An MPS can be encoded as the evolution of a product state (analogous to $\prod_{\otimes q=1}^{Q}\boldsymbol{v}$ in Eq. (2)) by implementing a quantum circuit formed by $O(1)$ layers of quantum gates (which is an unitary deep TN, analogous to our ADTN formed by several TN layers). Previous work showed that the number of parameters in a quantum circuit is generally much smaller than that in an MPS [40], suggesting that an MPS can be further compressed by deep TN. Our work was inspired by this observation, and experiments such as those in Fig. 3 support that ADTN’s with moderate depths ($O(1)$ TN layers) possess superior compression abilities compared to MPS or MPO.
IV Conclusion
In summary, we proposed a general compression scheme based on automatically differentiable tensor network (ADTN) to significantly reduce the variational parameters of neural networks (NN’s). The key idea is to encode the higher-order parameter tensors of NN as the contractions of deep ADTN(s) that contain exponentially fewer parameters. The performance of our scheme has been demonstrated with several well-known NN’s on different datasets. Discussions on overparametrization, compression order, faithfulness, and depth of ADTN were given. Our work suggests that deep tensor network offers a more efficient and compact mathematical representation of the variational parameters in NN’s compared to multi-way arrays and shallow tensor networks.
The high flexibility of the proposed ADTN scheme allows for further improvements in compression performance. Besides the tuning of hyperparameters (depth, bond dimensions, etc.) of ADTN, architectures other than the brick-wall type can be considered to construct the ADTN. The activation functions can be replaced by others such as leaky ReLU. Instead of compressing one NN layer with multiple ADTN’s, one might try to compress multiple NN layers into a single ADTN by concatenating their parameters into one tensor. This is a reflection of the generality of our scheme, and might be useful in special cases (e.g., when there exist redundant NN layers). Again, the types of NN layers make no difference to ADTN scheme.
Though ADTN performs impressively for compression, the decoding process (contraction) introduces additional computational cost to inference. This is because the contraction of ADTN and the inference of NN are currently treated as two separated procedures. One should first implement the contraction to get the parameters of NN, then followed by inference using the NN. Computational efficiency can be improved by combining these two procedures, which we leave for the future investigation.
@@ Acknowledgments
Y.Q. is thankful to Ze-feng Gao for stimulating discussions.
Funding: This work was supported in part by Beijing Natural Science Foundation (Grant No. 1232025), the Ministry of Education Key Laboratory of Quantum Physics and Photonic Quantum Information (Grant No. ZYGX2024K020), and Academy for Multidisciplinary Studies, Capital Normal University.
Author contributions: S.J.R. conceived and initiated the presented ideas, and supervised the research. Y.Q. and K.L. contributed to most of the coding works and numerical simulations. All authors contributed to the algorithms, analyses, and the preparation of the manuscript.
Competing interests: The authors declare that they have no competing interests.
Data availability
The data are available on request from the corresponding authors.
Appendix
Appendix A Necessity of pre-training
Fig. 6 compares the losses versus the training epochs during the fine-tuning stage, with or without pre-training. The results using the L1 and L2 (Euclidean) distances for pre-training are shown. In this experiment, we compress the last three layers of AlexNet simultaneously, excluding the output layer. Faster convergence and lower losses are achieved with pre-training. Meanwhile, the L2 distance results in a slightly lower loss than the L1 distance.
Figure 6: (Color online) Loss versus the training epochs in the second (fine-tuning) stage. The left panel shows the loss for ADTN(2-8) with three conditions: no pre-training, pre-training using L2 (Euclidean) distance, and pre-training using L1 distance. By “(2-8)”, we mean to take $2$ ADTN’s for compressing each NN layer, where each ADTN contains 8 tensor layers. The right panel presents the losses for ADTN(4-2).
Appendix B Hyperparameters of ADTN’s for the compressing VGG-16
In the whole work, we fix the dimensions of each tensor in the ADTN’s to be $(2\times 2\times 2\times 2)$, and take the brick-wall TN structure. The other key hyperparameters of the ADTN compression scheme are the number of ADTN’s, $M$, and the value of $Q$ for each ADTN (i.e., the order of tensor obtained by contracting the ADTN). Table 3 shows the values of $Q$ for the compression of the NN layers in VGG-16. We fix $M=2$. Each ADTN is configured with a brick-wall architecture, as shown in Fig. 7. For an NN layer, the number of parameters compressed by the two ADTN’s is given by $2^{Q_{1}}+2^{Q_{2}}$, where the values of $Q$ (denoted as $Q_{1}$ and $Q_{2}$ for the two ADTN’s, respectively) are taken as large and as equal as possible. Correspondingly, the number of parameters in the two ADTN’s is
$16M(Q_{1}+Q_{2}-2).$
Table 3: Sample of compressed VGG-16 with ADTN(N-M), where $N=2$, $M=4$.
Layer name Weight shape Params $Q_{1}$ $Q_{2}$ ADTN Params
Conv1 96$\times$3$\times$3$\times$3 - - - -
Conv2 96$\times$96$\times$3$\times$3 - - - -
Conv3 128$\times$96$\times$3$\times$3 - - - -
Conv4 128$\times$128$\times$3$\times$3 $2^{17}+2^{14}$ $17$ $14$ $1856$
Conv5 256$\times$128$\times$3$\times$3 $2^{18}+2^{15}$ $18$ $15$ $1984$
Conv6 256$\times$256$\times$3$\times$3 $2^{19}+2^{16}$ $19$ $16$ $2112$
Conv7 256$\times$256$\times$3$\times$3 $2^{19}+2^{16}$ $19$ $16$ $2112$
Conv8 512$\times$256$\times$3$\times$3 $2^{20}+2^{17}$ $20$ $17$ 2240
Conv9 512$\times$512$\times$3$\times$3 $2^{21}+2^{18}$ $21$ $18$ 2368
Conv10 512$\times$512$\times$3$\times$3 $2^{21}+2^{18}$ $21$ $18$ 2368
Conv11 512$\times$512$\times$3$\times$3 $2^{21}+2^{18}$ $21$ $18$ 2368
Conv12 512$\times$512$\times$3$\times$3 $2^{21}+2^{18}$ $21$ $18$ 2368
Conv13 512$\times$512$\times$3$\times$3 $2^{21}+2^{18}$ $21$ $18$ 2368
FC1 512$\times$4096 $2^{20}+2^{20}$ $20$ $20$ 2304
FC2 4096$\times$4096 $2^{23}+2^{23}$ $23$ $23$ 2688
FC3 4096$\times$×class - - - -
Appendix C Basics of tensor network and the contractions
@@
Appendix D Details of FC-2, LeNet-5, AlexNet, AFNet and VGG-16